OpenStudy (anonymous):

Solve for x Km/ 3 - 2x = 4 x=km-4 ----- 2 x=4-km ----- 6 x=2-km ----- 3 x=km-12 ----- 6

OpenStudy (anonymous):

km / 3 - 2x = 4 add 2x both side km/3=4+2x multy. by 3 km=12+6x 6x=km-12 x=(km-12)/6